OpenStudy (anonymous):

@ganeshie8 @jim_thompson5910

OpenStudy (anonymous):

What polynomial has roots of -6, -4, and 1? x^3 - 9x^2 - 22x + 24 x^3 - x^2 - 26x - 24 x^3 + x^2 - 26x + 24 x^3 + 9x^2 + 14x - 24

OpenStudy (anonymous):

i think its D

OpenStudy (anonymous):

I thinking D too

OpenStudy (anonymous):

thnx

Miracrown (miracrown):

Me three

OpenStudy (anonymous):

yay! i have a couple more do you mind helping me with ?

OpenStudy (anonymous):

Find the horizontal or oblique asymptote of f(x) =(-3x^2 + 7x + 1)/(x - 2) y = 2 y = -3 y = 3x + 7 y = -3x + 1

OpenStudy (anonymous):

i think this one is D too

Miracrown (miracrown):

Yes, you're right :) that is the oblique asymptote

OpenStudy (anonymous):

oh so i was righ t okay lol

Miracrown (miracrown):

yes lol

OpenStudy (anonymous):

i was getting confused

Miracrown (miracrown):

:)

OpenStudy (anonymous):

i got one more this one is easier but idk the aqnswer :(

OpenStudy (anonymous):

btw thnx

OpenStudy (anonymous):

Given a polynomial f(x), if (x + 2) is a factor, what else must be true? f(0) = 2 f(0) = -2 f(2) = 0 f(-2) = 0

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

i think its D not sure though or C but leaning more towarfs D

OpenStudy (unklerhaukus):

\[f(x) = (x+2)\cdot g(x)=0\]

OpenStudy (unklerhaukus):

when x = -2, .... you get D

OpenStudy (anonymous):

thnx your!!

OpenStudy (anonymous):

thank you ^* lol the other didn't make sense